# Gauge Symmetries, Contact Reduction, and Singular Field Theories

### v1

@@ IV Contact Geometry
In order to examine the commutative relationship between contact reduction by scaling symmetries, and the restriction of the phase space of mechanical systems described by singular Lagrangians, it will be necessary to understand how the geometric algorithm we have presented for pre-symplectic manifolds is to be adapted to the pre-contact spaces, which naturally arise upon eliminating a scaling degree of freedom [14, 15, 10]. In anticipation of this, we shall briefly introduce the fundamental ideas of contact and pre-contact geometry [59, 60], together with their implementation in the description of Lagrangian and Hamiltonian systems [61, 62, 63].
IV.1 Contact Manifolds
A contact structure on a smooth manifold $C$, of dimension $2n+1$, is a maximally-non-integrable distribution $\xi\subset TC$; locally, this distribution may be described as the kernel of some $\eta\in\Omega^{1}(U\subset C)$, referred to as a contact form. If the quotient line bundle $TC/\xi\rightarrow C$ is trivial, then $\xi$ is coorientable, and $\eta$ may be extended to the whole of $C$ [64]. The non-integrability condition of $\xi$ may be expressed as the requirement that
$\eta\wedge d\eta^{n}\neq 0$ (19)
over the whole of $C$. The coorientability of $\xi$ implies that $C$ itself is orientable; in this case, $\eta\wedge d\eta^{n}$ provides a volume form on $C$, and since, in general, $\xi$ is of rank $2n$, at each point $p\in C$, we have the decomposition
$T_{p}C=\textrm{ker}\,\eta_{p}\oplus\textrm{ker}\,d\eta_{p}$
and $d\eta_{p}$ defines a skew-symmetric, non-degenerate, bilinear form of maximal rank on the $2n$-dimensional subspace $\xi_{p}$. We henceforth assume that $\xi$ is coorientable, and work with the globally-defined form $\eta$; every contact manifold $(C,\eta)$ admits a distinguished Reeb vector field $\mathcal{R}\in\mathfrak{X}^{\infty}(C)$, defined via
@@ IV.1 Contact Manifolds
IV.2 Lagrangian Mechanics
Consider the $n$-dimensional configuration space $Q$, with local coordinates $(q^{i})$; the extended tangent bundle $TQ\times\mathbb{R}$ is a manifold of dimension $2n+1$, with induced bundle coordinates $(q^{i},v^{i},z)$. The canonical structures of $TQ$ are easily extended to $TQ\times\mathbb{R}$ [62], and have local coordinate expression (1). A contact Lagrangian is a function $L:TQ\times\mathbb{R}\rightarrow\mathbb{R}$, which allows us to define the Cartan forms $\theta_{L}$ and $\omega_{L}$, together with the energy function $E_{L}$. The space $TQ\times\mathbb{R}$ is made into a contact manifold upon introducing the 1-form $\eta_{L}:=dz-\theta_{L}$, and we refer to the triple $(TQ\times\mathbb{R},\eta_{L},E_{L})$ as a contact Lagrangian system.
IV.3 The Legendre Map
Whilst contact Lagrangian systems are introduced on the space $TQ\times\mathbb{R}$, their Hamiltonian counterparts are studied on the extended cotangent bundle $T^{*}Q\times\mathbb{R}$, which admits the following projections
$\begin{split}\pi_{0}:T^{*}Q\times\mathbb{R}\rightarrow Q\times\mathbb{R}&\quad\quad\quad\quad\pi_{1}:T^{*}Q\times\mathbb{R}\rightarrow T^{*}Q\\ \rho:\,&T^{*}Q\times\mathbb{R}\rightarrow\mathbb{R}\end{split}$ (21)
The Legendre map $FL:TQ\times\mathbb{R}\rightarrow T^{*}Q\times\mathbb{R}$ acts on the local bundle coordinates $(q^{i},p_{i},z)$ of $T^{*}Q\times\mathbb{R}$ according to
@@ IV.3 The Legendre Map
The extended cotangent bundle inherits a contact structure, defined in terms of the canonical structures of $T^{*}Q$; in particular, if $\vartheta\in\Omega^{1}(T^{*}Q)$ and $\Omega=-\,d\vartheta\in\Omega^{2}(T^{*}Q)$ denote the Liouville 1- and 2-forms on the cotangent bundle, then we have
$\theta=\pi_{1}^{*}\vartheta\quad\quad\textrm{and}\quad\quad\omega=-\,d\theta=\pi_{1}^{*}\Omega$ (23)
The Legendre map relates these forms with $\theta_{L}$ and $\omega_{L}$ in the obvious manner
$FL^{*}\theta=\theta_{L}\quad\quad\textrm{and}\quad\quad FL^{*}\omega=\omega_{L}$ (24)
The contact 1-form $\eta$ is then $\eta=dz-\theta$, and the Reeb vector field $\mathcal{R}$ follows from (20); in local coordinates, we have
$\theta=p_{i}\,dq^{i}\quad\quad\omega=dq^{i}\wedge dp_{i}\quad\quad\eta=dz-p_{i}\,dq^{i}\quad\quad\mathcal{R}=\frac{\partial}{\partial z}$ (25)
In complete analogy to the symplectic case, the Legendre map provides a set of criteria, according to which we may classify a given contact Lagrangian system; in particular, we shall restrict ourselves to those singular systems which are categorised as ‘almost regular.’
IV.4 Hamiltonian Formalism
If $(TQ\times\mathbb{R},\eta_{L},E_{L})$ is a hyperregular contact Lagrangian system, we introduce the Hamiltonian as the unique function $H:T^{*}Q\times\mathbb{R}\rightarrow\mathbb{R}$, such that $FL^{*}H=E_{L}$. The triple $(T^{*}Q\times\mathbb{R},\eta,H)$ is then said to constitute a contact Hamiltonian system, and there exists a bundle morphism
@@ IV.4 Hamiltonian Formalism
V Pre-contact Systems
All results presented thus far are applicable to regular and hyperregular systems; however, the main objective of this work is to present a mathematical framework within which to analyse scaling symmetries in singular theories. Consequently, we shall dedicate the following sections to developing a number of tools required to treat such systems.
V.1 Jacobi Structures
@@
V.2 Pre-contact Manifolds
The final piece of mathematical background required to introduce the constraint algorithm is that of pre-contact manifolds [62, 74]; suppose that $(M,\eta)$ is a $(2n+1)$-dimensional manifold, equipped with a 1-form $\eta$, such that the distribution
$\mathcal{C}:=\textrm{ker}\,\eta\,\cap\,\textrm{ker}\,d\eta\;\subseteq\;TM$ (38)
is of rank 2n+1p2n+1-p; in this case, we refer to $\mathcal{C}$ as the characteristic distribution, and say that $\eta$ is of class $p$. In general, if $\eta$ is of class $2k+1$, then we have
In addition, around each point $p\in M$, there exist local coordinates $(x_{1},\cdots,x_{k},y^{1},\cdots,y^{k},z,w_{1},\cdots,w_{s})$, with s=2(nk)s=2(n-k), such that $\eta$ is given by $\eta=dz-y_{i}\,dx^{i}$, and the characteristic distribution is spanned by
$\mathcal{C}\,=\,\,\biggr\langle\frac{\partial}{\partial w_{1}},\,\cdots,\frac{\partial}{\partial w_{s}}\biggr\rangle$
Upon introducing a function $H:M\rightarrow\mathbb{R}$, we refer to the triple $(M,\eta,H)$ as a pre-contact Hamiltonian system; as in the case of regular contact systems, we introduce the vector bundle morphism $\bar{\flat}:TM\rightarrow T^{*}M$, with precisely the same action as in (26).
Every pre-contact manifold admits a Reeb vector field $\mathcal{R}$, defined according to the conditions $\iota_{\mathcal{R}}\,d\eta=0$ and $\iota_{\mathcal{R}}\eta=1$; however, the novelty is that this field is no longer unique. In particular, any $X\in\mathfrak{X}^{\infty}(M)$ is a Reeb vector field if $\bar{\flat}(X)=\eta$ [60].
Given a distribution $\mathcal{D}\subset TM$, we define the following notion of orthogonality
$\mathcal{D}^{\perp}:=\{v\in TM\;|\;\bar{\flat}(w)(v)=0\quad\textrm{for all}\;w\in\mathcal{D}\,\}$ (40)
Note that since $\bar{\flat}(v)(w)$ is neither symmetric or antisymmetric under the exchange of $v$ and $w$, the above definition is strictly the right complement of $\mathcal{D}$; it is of course possible to define the left complement; however, $\mathcal{D}^{\perp}$ is the space we shall use within the constraint algorithm.
@@ VI Pre-contact Constraint Algorithm
We begin by considering a pre-contact Lagrangian system $(TQ\times\mathbb{R},\eta_{L},E_{L})$, of class $2k+1$, which occurs precisely when the rank of the Hessian matrix is equal to $k$ over the whole of $TQ\times\mathbb{R}$. The extended cotangent bundle $T^{*}Q\times\mathbb{R}$ has a canonical contact form $\eta$, whose construction was given in (23); if we introduce the inclusion $\kappa_{0}:P_{0}:=FL(TQ\times\mathbb{R})\hookrightarrow T^{*}Q\times\mathbb{R}$, then $P_{0}$ inherits the structure of a pre-contact manifold, with corresponding 1-form $\eta_{0}=\kappa_{0}^{*}\eta=FL_{0}^{*}\,\eta_{L}$, where $FL_{0}$ refers to the restriction of the Legendre map to $P_{0}$. This pre-contact manifold is then made into a Hamiltonian system upon introducing the unique function $H_{0}:P_{0}\rightarrow\mathbb{R}$, which satisfies $FL_{0}^{*}\,H_{0}=E_{L}$. In the interest of reducing cumbersome notation in what follows, we introduce
@@
and write the restriction of the bundle morphism $\bar{\flat}$ to $P_{0}$ as $\bar{\flat}_{0}$; thus, the dynamical problem we seek to solve is that of finding a vector field $X_{H}\in\mathfrak{X}^{\infty}(P_{0})$, such that
@@
As was the case for the pre-symplectic algorithm, it is convenient to make use of the relationship between the annihilator of the space $TP_{0}^{\perp}$ and the image of $TP_{0}$ under $\bar{\flat}_{0}$ to write $P_{1}$ as
$P_{1}=\{x\in P_{0}\;|\;\langle\,(\alpha_{0})_{x},T_{x}P_{0}^{\perp}\,\rangle=0\,\}$ (43)
Recall that for the pre-symplectic manifold $M_{0}$, we observed that $TM_{0}^{\perp}=\textrm{ker}\,\omega_{0}$, which led to the recharacterisation of $M_{1}$ given in (10); in the present context, we find that $TP_{0}^{\perp}=\mathcal{C}$, which provides the following, more practical condition for deducing $P_{1}$
@@
Upon finding a basis for the characteristic distribution, $P_{1}$ may be described as the zero-set of a collection of secondary constraint functions. We must now impose that any vector field solution on $P_{1}$ be tangent to this space, which generally requires us to restrict to the smaller set
@@
In order to deduce local expressions for the dynamical evolution of extended phase space functions, we shall suppose that $P_{f}$ may be written as the zero-set of the $N$ functions $\{\Phi^{\alpha}\}$. Introducing the matrix of brackets $K^{\alpha\beta}:=\{\Phi^{\alpha},\Phi^{\beta}\}_{J}$, with $\textrm{rank}\,K=N-F$, we extract $F$ first class functions $\Omega^{a}$, and NFN-F second class functions $\chi^{j}$, following precisely the same procedure as in (14).
@@
Upon expanding this expression, making use of the weak Leibnitz rule, we find that
$\dot{f}=\left(X_{H}+u_{a}X_{\Omega^{a}}\right)(f)+\Omega^{a}\Lambda_{DJ}(df,du_{a})$ (49)
When evaluated on the constraint surface $P_{f}$, the term proportional to $\Omega^{a}$ vanishes identically, leaving
@@ VII.1 Scaling Symmetries of Regular Hamiltonian Systems
Here, and throughout what follows, we use the symbol $\mathfrak{L}$ to denote the Lie derivative, reserving the calligraphic $\mathcal{L}$ for the Lagrangian density, encountered in our treatment of multisymplectic field theory.
As a consequence of (51), we see that if $X_{H}\in\mathfrak{X}^{\infty}(M)$ is the unique Hamiltonian vector field corresponding to $H$ via the relationship $\iota_{X_{H}}\omega=dH$, then
@@
Since $\omega$ is, by assumption, non-degenerate, it follows that $D,X_{H}]=(\Lambda-1)X_{H}$, so that the effect of $D$ is to rescale all trajectories by the same non-zero factor, and may thus be considered a generator of changes in global scale. The deduction of physical law is, fundamentally, an empirical process; a measurement of length, for instance, is only meaningful as compared to some fiducial length, taken to represent a ‘metre stick’ [77]. Under the effect of a scaling symmetry, not only is the length to be measured scaled, but so too is the instrument used to perform the measurement. Consequently, a dynamical similarity maps between configurations whose constituents are rescaled in such a way so as to produce no observable effect. Such a transformation is the mathematical realisation of Poincaré’s thought experiment, where one imagines waking up to a world in which all distances have suffered a doubling in scale [78]: any attempt to discern the rescaled world from the original configuration will be futile, since observers have access only to ratios of dimensionful quantities. We shall develop these ideas somewhat further in section (XII), in the context of low-energy effective gauge theories, examining arguments presented against Poincaré’s stance that a rescaled world would be indistinguishable from our own.
From a minimalistic viewpoint, a mathematical description whose ontology contains redundant degrees of freedom, pertaining to unobservable global scales, is highly undesirable [79, 80]. We argue that, where possible, any theory of the natural world should be formulated exclusively in terms of empirically accessible quantities, together with those (possibly unobservable) parameters, whose presence contributes to the closure of the algebra of dynamical observables. The identification of a scaling symmetry indicates that our mathematical framework contains just such a redundancy, and the role of ‘contact reduction’ is to make precise the heuristic notion of eliminating this superfluous structure.
@@
generates $\pi_{*}\,\textrm{span}(X_{H})$. In general, a contact vector field is one whose flow preserves the contact distribution; consequently, unless the scaling symmetry of interest is of degree one (which, throughout subsequent discussion, will often be the case) or we restrict ourselves to the zero-set of $H^{c}$, the vector field $X$ does not generally preserve $\xi=\textrm{ker}\,\eta$. Additionally, it is clear that this construction is dependent upon the relation $\mathfrak{L}_{D}X_{H}=(\Lambda-1)X_{H}$, which itself derives from the symplectic Hamiltonian equation of motion $\iota_{X_{H}}\omega=dH$; as a result, scaling symmetries are strictly dynamical in nature: that is to say, the contact-reduced system faithfully reproduces the original symplectic theory only when the equations of motion are satisfied.
@@
This concludes our introductory treatment of contact reduction, and we now apply these ideas to constrained Hamiltonian systems. Before commencing, we remark that there are two conceivable ways of proceeding: on the one hand, given a Hamiltonian $H_{0}$, corresponding to a singular Lagrangian, we may apply the techniques of section (III.1), deduce the final constraint manifold, eliminating any gauge degrees of freedom, and then seek a scaling symmetry, to make the contact reduction. Alternatively, we might first identify a scaling degree of freedom within the unconstrained Hamiltonian system, excise this from our ontology, and then use the pre-contact constraint algorithm, to deduce the subset of the reduced space, upon which solutions to the contact equations of motion are well-defined. More succinctly, in what follows, we would like to examine the commutativity of ‘restriction &\& reduction’ and ‘reduction &\& restriction.’
@@ VII.2 Reduction &\& Restriction
Let us suppose that $(TQ,\eta_{L},E_{L})$ is an almost-regular Lagrangian system; we begin by deducing the canonical Hamiltonian $H_{0}:M_{0}\rightarrow\mathbb{R}$, defined on the primary constraint manifold $\jmath_{0}:M_{0}\hookrightarrow T^{*}Q$. This is then extended to a function $H:T^{*}Q\rightarrow\mathbb{R}$ on the full phase space, such that $H|_{M_{0}}=H_{0}$. [^4] In practice, $H$ and $H_{0}$ have identical coordinate expressions; however, formally, they belong to different spaces. The cotangent bundle $T^{*}Q$ is a symplectic manifold, whose corresponding 2-form is expressed in local Darboux coordinates as $\omega=dq^{i}\wedge dp_{i}$; supposing that $M_{0}$ is described as the zero-set of the primary constraint functions $\phi^{\alpha}$, we seek a vector field $D\in\mathfrak{X}^{\infty}(T^{*}Q)$ which satisfies
@@
We highlight that this construction does not directly give rise to a pre-contact manifold; instead, we have symmetry reduced the full phase space, into which the pre-symplectic system is embedded. Had we sought a vector field on $M_{0}$ as our scaling symmetry, it is not clear that we should always obtain a well-defined pre-contact form on the quotient space. By working on the full phase space, we have the map $\beta:T^{*}Q\rightarrow C$, which allows us to construct functions $\gamma^{\alpha}\in C^{\infty}(C)$, whose zero-set defines a submanifold $\kappa_{0}:C_{0}\hookrightarrow C$, which does inherit a pre-contact structure; in particular, we have

### v2

@@ IV Contact Geometry
Contact structures arise naturally when scaling degrees of freedom are eliminated from the ontology of theories describing mechanical systems [14, 15, 10]. In general, a contact structure on a smooth manifold $C$, of dimension $2n+1$, is a maximally-non-integrable distribution $\xi\subset TC$; locally, this distribution may be described as the kernel of some $\eta\in\Omega^{1}(U\subset C)$, referred to as a contact form [49, 50, 51, 52, 53]. If the quotient line bundle $TC/\xi\rightarrow C$ is trivial, as we shall assume to be the case, then $\xi$ is coorientable, and $\eta$ may be extended to the whole of $C$ [54]. Every contact manifold $(C,\eta)$ admits a distinguished Reeb vector field $\mathcal{R}\in\mathfrak{X}^{\infty}(C)$, defined via
@@
The extended tangent bundle $TQ\times\mathbb{R}$ of the $n$-dimensional configuration space $Q$ is a manifold of dimension $2n+1$, with local coordinates $(q^{i},v^{i},z)$. A contact Lagrangian is a function $L:TQ\times\mathbb{R}\rightarrow\mathbb{R}$, from which we define the Cartan forms $\theta_{L}$ and $\omega_{L}$, together with the energy function $E_{L}$, exactly as on $TQ$. The space $TQ\times\mathbb{R}$ is made into a (pre-)contact manifold, introducing the 1-form $\eta_{L}:=dz-\theta_{L}$, and we refer to the triple $(TQ\times\mathbb{R},\eta_{L},E_{L})$ as a (pre-)contact Lagrangian system.
If the characteristic distribution
$\mathcal{C}:=\textrm{ker}\,\eta_{L}\,\cap\,\textrm{ker}\,d\eta_{L}\;\subseteq\;T(TQ\times\mathbb{R})$ (16)
is of rank 2(nk)2(n-k), we say that $\eta_{L}$ is of class $2k+1$ [52, 55]; locally, this implies
Clearly, when $\textrm{rank}\,\mathcal{C}=2n$, $(TQ\times\mathbb{R},\eta_{L},E_{L})$ is a regular system; in such a case, the Legendre map $FL:TQ\times\mathbb{R}\rightarrow T^{*}Q\times\mathbb{R}$ is a diffeomorphism, and acts on local coordinates $(q^{i},p_{i},z)$ of $T^{*}Q\times\mathbb{R}$ according to
@@
The space $T^{*}Q\times\mathbb{R}$ has a canonical contact form $\eta=dz-p_{i}\,dq^{i}$, and the Reeb vector field $\mathcal{R}$ follows from (15).
If $(TQ\times\mathbb{R},\eta_{L},E_{L})$ is a hyperregular Lagrangian system, we introduce the unique function $H:T^{*}Q\times\mathbb{R}\rightarrow\mathbb{R}$, such that $FL^{*}H=E_{L}$. The triple $(T^{*}Q\times\mathbb{R},\eta,H)$ is then said to constitute a contact Hamiltonian system, and there exists a bundle morphism
@@
V Jacobi Structures
@@ VI Pre-contact Constraint Algorithm
When $(TQ\times\mathbb{R},\eta_{L},E_{L})$ is a pre-contact system, the above construction requires only minor modification. The image $P_{0}:=FL(TQ\times\mathbb{R})$ is assumed to be a closed submanifold of $T^{*}Q\times\mathbb{R}$, with inclusion $\kappa_{0}:P_{0}\hookrightarrow T^{*}Q\times\mathbb{R}$; the canonical contact form on $T^{*}Q\times\mathbb{R}$ restricts to $\eta_{0}=\kappa_{0}^{*}\eta$, and endows $P_{0}$ with a pre-contact structure. This is then made into a Hamiltonian system upon introducing the unique function $H_{0}:P_{0}\rightarrow\mathbb{R}$, which satisfies $FL_{0}^{*}\,H_{0}=E_{L}$, where $FL_{0}$ denotes the restriction of $FL$ to its image. Denoting the restriction of the bundle morphism $\bar{\flat}$ to $P_{0}$ via $\bar{\flat}_{0}$, for a distribution $\mathcal{D}\subset TP_{0}$, we have the following notion of orthogonality
$\mathcal{D}^{\perp}:=\{v\in TP_{0}\;|\;\bar{\flat}_{0}(w)(v)=0\quad\textrm{for all}\;w\in\mathcal{D}\,\}$ (31)
In the interest of reducing cumbersome notation in what follows, we introduce
@@
so that the dynamical problem to be solved is that of finding a vector field $X_{H}\in\mathfrak{X}^{\infty}(P_{0})$, such that
@@
As was the case for the pre-symplectic algorithm, it is convenient to make use of the relationship $(TP_{0}^{\perp})^{o}=\bar{\flat}_{0}(TP_{0})$; further, it is found that $TP_{0}^{\perp}=\mathcal{C}$, which provides the following, more practical condition for deducing $P_{1}$
@@
We must now impose that any vector field solution on $P_{1}$ be tangent to this space, which generally requires us to restrict to the smaller set
@@
In order to present the complementary local description, we shall suppose that $P_{f}$ may be written as the zero-set of the $N$ functions $\{\Phi^{\alpha}\}$. Introducing the matrix of brackets $K^{\alpha\beta}:=\{\Phi^{\alpha},\Phi^{\beta}\}_{J}$, with $\textrm{rank}\,K=N-F$, we extract $F$ first class functions $\Omega^{a}$, and NFN-F second class functions $\chi^{j}$, following precisely the same procedure as in (10).
@@
When evaluated on the constraint surface $P_{f}$, the constraints $\Omega^{a}$ vanish identically, and we find
@@ VII.1 Scaling Symmetries of Regular Hamiltonian Systems
Here, $\mathfrak{L}$ denotes the Lie derivative; we reserve the calligraphic $\mathcal{L}$ for the Lagrangian density. As a consequence of (40), we see that if $X_{H}\in\mathfrak{X}^{\infty}(M)$ is the unique Hamiltonian vector field corresponding to $H$ via the relationship $\iota_{X_{H}}\omega=dH$, then
@@
Since $\omega$ is, by assumption, non-degenerate, it follows that $D,X_{H}]=(\Lambda-1)X_{H}$, so that the effect of $D$ is to rescale all trajectories by the same non-zero factor, and may thus be considered a generator of changes in global scale. The deduction of physical law is, fundamentally, an empirical process; a measurement of length, for instance, is only meaningful as compared to some fiducial length, taken to represent a ‘metre stick’ [67]. Under the effect of a scaling symmetry, not only is the length to be measured scaled, but so too is the instrument used to perform the measurement. Consequently, a dynamical similarity maps between configurations whose constituents are rescaled in such a way so as to produce no observable effect. Such a transformation is the mathematical realisation of Poincaré’s thought experiment, where one imagines waking up to a world in which all distances have suffered a doubling in scale [68]: any attempt to discern the rescaled world from the original configuration will be futile, since observers have access only to ratios of dimensionful quantities.
From a minimalistic viewpoint, a mathematical description whose ontology contains redundant degrees of freedom is highly undesirable [69, 70]. We argue that, where possible, any theory of the natural world should be formulated exclusively in terms of empirically accessible quantities, together with those (possibly unobservable) parameters, whose presence contributes to the closure of the algebra of dynamical observables. The identification of a scaling symmetry indicates that our mathematical framework contains just such a redundancy, and the role of ‘contact reduction’ is to make precise the notion of eliminating this superfluous structure.
@@
generates $\pi_{*}\,\textrm{span}(X_{H})$. A contact vector field is one whose flow preserves the contact distribution; consequently, unless the scaling symmetry of interest is of degree one (which will often be the case) or we restrict ourselves to the zero-set of $H^{c}$, the vector field $X$ does not generally preserve $\xi=\textrm{ker}\,\eta$. Additionally, it is clear that this construction is dependent upon the relation $\mathfrak{L}_{D}X_{H}=(\Lambda-1)X_{H}$, which itself derives from the symplectic Hamiltonian equation of motion $\iota_{X_{H}}\omega=dH$; as a result, scaling symmetries are strictly dynamical in nature: that is to say, the contact-reduced system faithfully reproduces the original symplectic theory only when the equations of motion are satisfied.
@@
This concludes our introductory treatment of contact reduction, and we now apply these ideas to constrained Hamiltonian systems. Note that there are two conceivable ways of proceeding: on the one hand, given a Hamiltonian $H_{0}$, corresponding to a singular Lagrangian, we may apply the techniques of section (III.1), deduce the final constraint manifold, and then make the contact reduction. Alternatively, we might first identify a scaling degree of freedom within the unconstrained Hamiltonian system, excise this from our ontology, and then use the pre-contact constraint algorithm, to deduce the subset of the reduced space upon which solutions to the contact equations of motion are well-defined.
@@ VII.2 Reduction &\& Restriction
Suppose that $(TQ,\eta_{L},E_{L})$ is an almost-regular Lagrangian system; we begin by deducing the canonical Hamiltonian $H_{0}:M_{0}\rightarrow\mathbb{R}$, defined on the primary constraint manifold $\jmath_{0}:M_{0}\hookrightarrow T^{*}Q$. This is then extended to a function $H:T^{*}Q\rightarrow\mathbb{R}$ on the full phase space, such that $H|_{M_{0}}=H_{0}$. [^3] In practice, $H$ and $H_{0}$ have identical coordinate expressions; however, formally, they belong to different spaces. The cotangent bundle $T^{*}Q$ is a symplectic manifold, whose corresponding 2-form is expressed in local Darboux coordinates as $\omega=dq^{i}\wedge dp_{i}$; supposing that $M_{0}$ is described as the zero-set of the primary constraint functions $\phi^{\alpha}$, we seek a vector field $D\in\mathfrak{X}^{\infty}(T^{*}Q)$ which satisfies
@@
We highlight that this construction does not directly give rise to a pre-contact manifold; instead, we have symmetry reduced the full phase space, into which the pre-symplectic system is embedded. The map $\beta:T^{*}Q\rightarrow C$ allows us to construct functions $\gamma^{\alpha}\in C^{\infty}(C)$, whose zero-set defines a submanifold $\kappa_{0}:C_{0}\hookrightarrow C$, which does inherit a pre-contact structure; in particular, we have